OpenStudy (pphalke):

A coin collector has $1.85 in dimes and nickels. she has 8 more dimes than nickels. how many nickels does she have?

OpenStudy (kropot72):

Let n be the number of nickels. Now we can write the equation: 5n + 10(n + 8) = 185 Next you need to solve this equation to find the value of n.

OpenStudy (pphalke):

5n+10n+80=185? is this right once I solved

OpenStudy (pphalke):

kropot? am I right

OpenStudy (kropot72):

Yes, you are right so far.

OpenStudy (pphalke):

15n+80=185 is this right once I collect like terms

OpenStudy (kropot72):

Yes, you are still right.

OpenStudy (pphalke):

so now I just subtract 80 from both sides

OpenStudy (kropot72):

Yes, that is the next step.

OpenStudy (pphalke):

15n=105?

OpenStudy (kropot72):

Yes, good work. Now divide both sides by 15.

OpenStudy (pphalke):

7 I get that diving 105/15

OpenStudy (kropot72):

15n/15 = n = 105/15 = 7 So, the number of nickels is 7.

OpenStudy (kropot72):

@pphalke Do you understand the last step?

OpenStudy (pphalke):

yes I did